The proposed Sky train fare would charge ₱35for the first 5 km of travel and ₱6.50for each additional kilometer over the proposed fare.
A. Find the fare function f(s) where s represents the number of kilometers needed.
B. What is the proposed fare for a distance of 50km?
C. How much isthe fare for the distance of 45km?
D. Jay will travel to UP Diliman, he travelled a distance of 23km. How much did he pay for the train fare?

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

The proposed Sky train fare would charge ₱35for the first 5 km of travel and ₱6.50for each additional kilometer over the proposed fare.
:
A. Find the fare function f(s) where s represents the number of kilometers needed.
f(s) = 6.5(s-5) + 35
:
B. What is the proposed fare for a distance of 50km?
f(s) = 6.5(50-5) + 35
f(s) = 6.5(45) + 35
f(s) = 292.5 + 35
f(s) = P327.50
:
C. How much is the fare for the distance of 45km?
Do this the same way, s=45
:
D. Jay will travel to UP Diliman, he travelled a distance of 23km.
How much did he pay for the train fare?
s = 23

Answer by ikleyn(52812)

.
The proposed Sky train fare would charge ₱35for the first 5 km of travel and ₱6.50for each additional kilometer over the proposed fare.
:
A. Find the fare function f(s) where s represents the number of kilometers needed.
~~~~~~~~~~~~~~~~~~~~~~


For a beginner student/reader,  it is very important to know and to learn
on how to write such functions (piece-wise functions) correctly.

In this problem,  the definition of the function  f(s)  is a standard TWO-LINE form


             /  35,                 if  s <= 5 kilometers
            |
           /
   f(s) = <
           \
            |
             \  35 + 6.50*(s-5),    if  s > 5 kilometers.